# PO-Flow: Flow-based Generative Models for Sampling Potential Outcomes and Counterfactuals

### v1

@@ Abstract
We propose PO-Flow, a novel continuous normalizing flow (CNF) framework for causal inference that jointly models potential outcomes and counterfactuals. Trained via flow matching, PO-Flow provides a unified framework for individualized potential outcome prediction, counterfactual predictions, and uncertainty-aware density learning. Among generative models, it is the first to enable density learning of potential outcomes without requiring explicit distributional assumptions (e.g., Gaussian mixtures), while also supporting counterfactual prediction conditioned on factual outcomes in general observational datasets. On benchmarks such as ACIC, IHDP, and IBM, it consistently outperforms prior methods across a range of causal inference tasks. Beyond that, PO-Flow succeeds in high-dimensional settings, including counterfactual image generation, demonstrating its broad applicability.
@@ 1 Introduction
Predicting potential outcomes (POs) lies at the core of causal inference. In a typical setting, each individual has two potential outcomes: $Y^{(1)}$ if they receive a treatment ($A=1$) and $Y^{(0)}$ if they do not ($A=0$). However, in the real world, only one outcome is observed, while the other potential outcome remains fundamentally unknown [28]. Accurately estimating both POs is crucial for decision-making in domains such as healthcare [41, 27, 50], where clinicians may wish to compare a patient’s likely outcomes under different treatment options [60, 20].
@@
In the causal inference literature, the vast majority of methods focus on estimating the average treatment effect (ATE) and the conditional average treatment effect (CATE) at the population level [32, 25, 1, 22, 37, 42, 55, 67]. It is important to note that although some methods perform well in estimating population-level CATE, they may underperform in individualized potential outcome (PO) prediction, as accurately predicting POs in finite samples is challenging [52].
@@
However, making decisions for individuals based on average treatment effects can be misleading. Therefore, accurately predicting individualized POs and learning their full distributions with uncertainty is a crucial task [69, 35]. Recent works [67, 44, 45] have explored generative models to sample POs and estimate empirical distributions. However, methods that learn the full density of POs without relying on explicit distributional assumptions (e.g., Gaussian mixtures) remain scarce.
@@
In addition, another critical challenge is estimating the counterfactual outcome conditioned on the factual—that is, predicting what would have happened under the alternative treatment, given the observed outcome. This task is especially relevant in post-hoc clinical analyses, where practitioners ask, “Would the patient have fared better on another therapy?” [66, 29, 21].
@@
Counterfactual prediction differs from standard PO estimation—it requires reasoning about outcomes under an alternative treatment conditioned on the observed outcome, introducing dependencies between observed and unobserved outcomes. Structural Causal Models (SCMs) [48] and modern extensions [46, 6, 13, 7, 65, 49, 9] offer the most widely used theoretical framework for counterfactual reasoning. SCMs methods typically assume known causal structures or aim to learn them explicitly. Yet, it may be possible to estimate counterfactuals directly without recovering the SCM [70]. In PO-Flow, we do not explicitly assume an SCM, but seek to directly estimate counterfactual outcomes.
We aim to develop a generative modeling framework that jointly supports population-level CATE estimation, individualized PO prediction with density modeling, and counterfactual outcome prediction without explicitly modeling an SCM—an area that has been relatively less explored.
@@ 1.1 Related Works
CATE Estimation: Numerous methods have been proposed for CATE estimation [32, 25, 1, 22, 37, 42, 55, 67]. Broadly, these fall into model-agnostic and model-specific approaches. Model-agnostic estimators, or meta-learners [37, 15, 16, 2, 54], typically follow one of two strategies [43]: (a) one-step methods, which fit regression models and compute CATE as the difference in predicted outcomes, and (b) two-step methods, which estimate nuisance functions to construct pseudo-outcomes, then regress them on covariates.
@@
Individualized POs Prediction: Although some CATE methods can be adapted for potential outcomes (POs) prediction, they may perform suboptimally, particularly on finite samples, as they are not specifically optimized for accurate PO estimation. This limitation partly arises because many methods assume that CATE has a simpler functional structure than POs. Nevertheless, several CATE approaches are also designed with POs prediction as the target [37, 55, 67, 42, 45, 43], and these will serve as our primary baselines for experimental comparisons.
@@
While SCMs offer theory-grounded counterfactual reasoning, general-purpose methods for individualized counterfactual prediction—without explicitly modeling SCMs or relying on auxiliary information—remain scarce. Although [67, 43] were not originally designed for counterfactuals, we adapt them by incorporating factual information into intermediate steps and use them as baselines.
Diffusion and Flow Models: Diffusion models [26, 57, 58] and flow-based generative models [12, 64, 40, 4] have demonstrated strong performance across a range of applications, including image generation, statistical sampling [18, 59, 61], and physics-informed modeling [8, 68, 51].
Recently, diffusion and flow-based generative models have been applied to causal inference tasks such as counterfactual image generation [53, 19], causal query answering [10], and inference over causal graphs or structural causal models [9, 33].
However, to our knowledge, only two generative algorithms have been proposed specifically for the potential outcomes (PO) task: the diffusion-based DiffPO [43] and the discrete normalizing flow-based INFs [45]. We conduct extensive comparisons with both methods in this work.
1.2 Main Contributions
We propose PO-Flow, a continuous normalizing flow (CNF) framework for potential outcomes and counterfactuals. Our main contributions are: (1) PO-Flow provides a unified generative framework for potential outcome prediction, CATE estimation, and general-purpose counterfactual prediction—without explicitly modeling structural causal models (SCMs) or relying on auxiliary information for the counterfactual task; (2) PO-Flow enables learning the density of predicted potential outcomes, offering deeper insights into uncertainty; (3) PO-Flow outperforms state-of-the-art methods across a range of causal tasks on multiple benchmarks.
@@ 2.1 Neural ODE and Continuous Normalizing Flow:
$\frac{dy(t)}{dt}=v_{\theta}(y(t),t),\quad t\in[0,1],$ (1)
where $v_{\theta}:\mathbb{R}^{d_{Y}+1}\to\mathbb{R}^{d_{Y}}$ represents the velocity field, which is parameterized by a neural network with parameter $\theta$. The time horizon is rescaled to $[0,1]$ without loss of generality.
@@
$\partial_{t}p(y,t)+\nabla\cdot(p(y,t)v_{\theta}(y,t))=0,\quad t\in[0,1],$ (2)
[…]
When the Neural ODE is well-posed, it defines a continuous and invertible mapping from the initial sample $y_{0}$ to the terminal sample $y_{1}$. The inverse mapping from $y_{1}$ to $y_{0}$ can be computed by integrating the ODE (1) in reverse time. This framework allows setting $y_{0}\sim p(\cdot,0)$ as the data distribution and $y_{1}\sim p(\cdot,1)$ as the base (noise) distribution, typically chosen as $N(0,I)$.
For notational convenience, we define $p_{Y}(\cdot):=p(\cdot,0)$ as the data distribution and $q(\cdot):=p(\cdot,1)$ as the base (noise) distribution. Throughout this paper, we adopt the common choice $q(\cdot)=N(0,I)$.
@@ 2.2 Flow Matching
$\mathcal{L}_{\text{FM}}(\theta)=\mathbb{E}_{t\sim\mathcal{U}[0,1],y\sim p(\cdot,t)}\left[\|v_{\theta}(y,t)-u(y,t)\|^{2}\right],$ (3)
@@
$\phi_{t}=(1-t)y_{0}+(t+\sigma_{\text{min}}(1-t))y_{1},$ (5)
where $y_{1}\sim N(0,I)$ and $\sigma_{\text{min}}$ is a small positive hyperparameter that ensures $p(\phi_{t},0)\sim N(y_{0},\sigma_{\text{min}}^{2}I)$. When $\sigma_{\text{min}}=0$, this reduces to the strict linear interpolant $\phi_{t}=(1-t)y_{0}+ty_{1}$.
Here, $\phi_{t}$ serves as a pre-specified reference trajectory that the ODE solution $y(t)$ should follow at time $t$. Under the linear interpolant defined in (5), the corresponding reference velocity field becomes:
$\frac{d\phi_{t}}{dt}=(1-\sigma_{\text{min}})y_{1}-y_{0}.$ (6)
@@
$\mathcal{L}_{\text{CFM}}(\theta)=\ \mathbb{E}_{t\sim\mathcal{U}[0,1],y_{0}\sim p_{Y}(\cdot),y_{1}\sim q(\cdot)}\|v_{\theta}(\phi_{t},t)-\frac{d\phi_{t}}{dt}\|^{2}.$ (7)
@@ 3 Problem Settings
Notations and Setups: We consider a setting where each individual is described by covariates $X\in\mathbb{R}^{d_{X}}$, a treatment assignment $A\in\{0,1\}$, and an outcome $Y^{(A)}\in\mathbb{R}^{d_{Y}}$ under treatment $A$. The observational dataset is $\mathcal{D}=\{(y^{(a)}_{i},x_{i},a_{i})\}_{i=1}^{n}\sim p_{Y,X,A}(\cdot)$, where $y_{i}^{(a)}$ is the observed potential outcome under treatment $a_{i}$, and $(x_{i},a_{i})\sim p_{X,A}(\cdot)$ are the associated covariates and treatment.
@@
Counterfactual Outcomes: In contrast to potential outcomes, we define the counterfactual outcome as $Y_{\text{cf}}^{(1-a)}:=Y^{(1-a)}\mid X,A=a,Y^{(a)}$: the outcome had the individual received the opposite treatment, conditioned on its observed potential (factual) outcome. For convenience, we denote $Y_{\text{cf}}^{(1-a)}$ and $\hat{Y}_{\text{cf}}^{(1-a)}$ as the true and predicted counterfactual outcomes, respectively.
To ensure the identifiability of potential outcomes, we adopt the standard assumptions commonly used in the literature [43, 45, 67]:
@@
(1) Consistency: If an individual receives treatment $A=a$, the observed outcome equals the corresponding potential outcome: $Y=Y^{(a)}$. (2) Unconfoundedness: All confounders are observed, i.e., $\{Y^{(0)},Y^{(1)}\}\perp\!\!\!\perp A\mid X$. (3) Overlap: Treatment assignment is probabilistic: $10<P(A=1\mid X=x)<$ 1, for all $x\in\mathcal{X}$.
Counterfactuals are rarely identifiable without strong assumptions [63, 47]. Estimating them can often be framed as an optimal transport problem [11, 17]. To constrain the space of joint transport mappings, our PO-Flow framework introduces the following assumption, applicable only to Section 4.2:
Assumption 3.2.
There exists a deterministic and invertible mapping $f:\mathbb{R}^{d_{Y}+d_{X}}\times\{0,1\}\rightarrow\mathbb{R}^{d_{Z}}$ that encodes the observed outcome as $Z_{Y}=f(Y^{(A)},X,A)$. A counterfactual latent representation can be obtained as $Z^{\prime}_{Y}=Z_{Y}+\delta$, where $\delta\sim\mathcal{N}(0,\Sigma(X))$ is an optional covariate-dependent noise term. The counterfactual is then recovered via $Y_{\text{cf}}^{(1-A)}=f^{-1}(Z^{\prime}_{Y},X,1-A)$.
@@ 4 PO-Flow: A Continuous Normalizing Flow for Potential Outcomes
We introduce PO-Flow, a continuous normalizing flow framework for modeling potential and counterfactual outcomes, with three main goals: 1. predicting individualized potential outcomes, 2. estimating counterfactual outcomes, and 3. learning the densities.
Our PO-Flow framework is built on Continuous Normalizing Flows (CNFs) and is trained using the Flow Matching (FM) objective. Our learning target is the trajectory of the sample $y(t)$ over $t\in[0,1]$. To condition the dynamics on both covariates $x$ and treatment $a$, our Neural ODE is defined as:
$\small\frac{dy(t)}{dt}=v_{\theta}(y(t),t,x,a),\quad t\in[0,1],$ (8)
where $v_{\theta}:\mathbb{R}^{d_{Y}+d_{X}+2}\rightarrow\mathbb{R}^{d_{Y}}$ is parametrized by a neural network.
@@
$\min_{\theta}\ \mathcal{L}_{\text{CFM}}(\theta)=\ \mathbb{E}_{t\sim\mathcal{U}[0,1],p_{Y,X,A}(\cdot),q(\cdot)}\|v_{\theta}(\phi_{t},t,x,a)-\frac{d\phi_{t}}{dt}\|^{2},$ (9)
where $\phi_{t}$ and $\frac{d\phi_{t}}{dt}$ are defined in (5) and (6), respectively.
Remark: Following prior work [23, 43], one may estimate propensity scores $w(x)=p(A=1\mid X=x)$ using a pre-trained classifier and apply inverse probability weighting $\frac{1}{w(x)}$ in the loss. However, we found this did not improve performance, but both implementations are included in our code.
Forward Process: Given the velocity field $v_{\theta}$, the forward process refers to pushing a factual outcome $y^{(a)}$ (at $t=0$) to a latent embedding $z_{y}$ (at $t=1$) via:
[…]
Through the forward process, $z_{y}$ encodes information from the factual outcome $y^{(a)}$ as well as its associated covariates $x$ and treatment $a$.
Reverse Process: The reverse process is initialized differently for interventional and counterfactual outcome prediction, as detailed in Sections 4.1 and 4.2. In general, given a noise sample $y_{1}\sim q(\cdot)$ at $t=1$, the reverse process is defined as:
[…]
Empirically, we use the Runge-Kutta method for numerical integration, with implementation details provided in Appendix B.4.
During training, we construct synthetic paths $\phi_{t}$ between $y_{0}$ and $y_{1}\sim q(\cdot)$ using the CFM loss, conditioning on $y_{0}$. However, Proposition 4.1 shows that the gradient of the CFM loss with respect to the model parameters is equivalent to the gradient of the original FM loss.
Proposition 4.1.
Assume that $p_{t}(y|x,a)>0$ for all $y\in\mathbb{R}^{d_{Y}}$ and $t\in[0,1]$. Then, up to a constant independent of $\theta$, the Conditional Flow Matching (CFM) loss and the original Flow Matching (FM) loss are equivalent. Hence,
As a result of Proposition 4.1, the learned velocity field $v_{\theta}$ during the reverse process generates samples $\hat{y}_{0}$ independent of the individual samples $y_{0}$ used during training. Consequently, since $y_{1}\sim q(\cdot)$ is random, $v_{\theta}$ induces a distribution over outcomes conditioned on $(x,a)$.

### v2

@@ Abstract
Predicting potential and counterfactual outcomes from observational data is central to clinical decision-making, where physicians must weigh treatments for an individual patient rather than relying solely on average effects at the population level. We propose PO-Flow, a continuous normalizing flow (CNF) framework for causal inference that jointly models potential outcomes and counterfactuals. Trained via flow matching, PO-Flow provides a unified approach to average treatment effect estimation, individualized potential outcome prediction, and counterfactual prediction. Besides, PO-Flow directly learns the densities of potential outcomes, enabling likelihood-based evaluation of predictions. Furthermore, PO-Flow explores counterfactual outcome generation conditioned on the observed factual in general observational datasets, with a supporting recovery result under certain assumptions. PO-Flow outperforms modern baselines across diverse datasets and causal tasks in the potential outcomes framework.
@@ 1 Introduction
Predicting potential outcomes (POs) lies at the core of causal inference. In a typical setting, each individual has two potential outcomes: $Y^{(1)}$ if they receive a treatment ($A=1$) and $Y^{(0)}$ if they do not ($A=0$). However, in the real world, only one outcome is observed, while the other potential outcome remains fundamentally unknown [25]. Accurately estimating both POs is crucial for decision-making in healthcare [37, 24, 46], where clinicians may wish to compare a patient’s likely outcomes under different treatment options rather than rely solely on average effects [56, 17].
@@
In the Potential Outcomes (POs) literature, the vast majority of methods focus on estimating the conditional average treatment effect (CATE): $\mathbb{E}[Y^{(1)}-Y^{(0)}|X]$, at the population level [28, 22, 1, 19, 33, 38, 50, 61]. It is important to note that although some methods perform well in estimating population-level CATE, they may underperform in individualized POs prediction, as accurately predicting POs in finite samples is challenging [47].
@@
However, making decisions for individuals based on average treatment effects can be misleading. Therefore, accurately predicting individualized POs and learning their full distributions $p(Y^{(A)}|X,A)$ is a crucial task [63, 31]. Recent works [61, 40, 41] have explored generative models to sample POs and estimate empirical distributions. However, methods that learn the full density of POs without relying on explicit distributional assumptions (e.g., Gaussian mixtures) remain scarce.
@@
In addition, another critical challenge is estimating the counterfactual outcome conditioned on the factual: $Y_{\mathrm{CF}}^{(1-A)}:=Y^{(1-A)}|X,Y^{(A)}$, i.e., predicting what would have happened under the alternative treatment, given the observed outcome $Y^{(A)}$. This task is especially relevant in post-hoc clinical analyses, where practitioners ask, “Would the patient have fared better on another therapy?” [60, 18].
@@
Counterfactual prediction differs from standard PO estimation—it requires reasoning about outcomes under an alternative treatment conditioned on the observed factual outcome, introducing dependencies between observed and unobserved outcomes. Structural Causal Models (SCMs) [44] and modern extensions [42, 6, 12, 7, 59, 45, 8] offer the most widely used theoretical framework for counterfactual reasoning. SCMs methods typically assume known causal structures or aim to learn them explicitly. However, it may be possible to estimate counterfactuals directly without recovering the SCM [64]. In PO-Flow, we do not explicitly assume an SCM, but seek to directly estimate counterfactual outcomes.
Despite recent progress, two key gaps remain in the potential outcomes literature. First, counterfactual outcome estimation conditioned on the observed factual outcome has received little attention, though it is crucial for post-hoc analyses. Second, beyond counterfactuals, most methods are specialized—some target CATE estimation, others aim at potential outcome prediction or distributional modeling—but few provide a unified and scalable algorithmic framework that can jointly address these tasks.
1.1 Main Contributions
We propose PO-Flow, a continuous normalizing flow (CNF) framework for potential outcomes and counterfactuals. Our main contributions are: (1) Enable counterfactual outcome generation conditioned on the observed factual, together with a supporting counterfactual recovery result (Corollary 4.5) under certain assumptions; (2) Provide a unified generative framework for potential outcome prediction, CATE estimation, and general-purpose counterfactual prediction; (3) Learn the density of potential outcomes, supporting principled uncertainty quantification; (4) Achieve superior performance over state-of-the-art methods across diverse causal tasks and benchmark datasets.
@@ 1.2 Related Works
CATE Estimation: Numerous methods have been proposed for CATE estimation [28, 22, 1, 19, 33, 38, 50, 61]. Broadly, these fall into model-agnostic and model-specific approaches. Model-agnostic estimators, or meta-learners [33, 15, 14, 2, 49], typically follow two strategies [39]: (a) one-step methods, which fit regression models and compute CATE as the difference in predicted outcomes, and (b) two-step methods, which estimate nuisance functions to construct pseudo-outcomes, then regress them on covariates.
@@
Individualized POs Prediction: Although some CATE methods can be adapted for potential outcomes (POs) prediction, they may perform suboptimally, particularly on finite samples, as they are not specifically optimized for accurate PO estimation. This limitation partly arises because many methods assume that CATE has a simpler functional structure than POs. Nevertheless, several CATE approaches are also designed with POs prediction as the target [33, 50, 61, 38, 41, 39].
@@
While SCMs offer theory-grounded counterfactual reasoning, general-purpose methods for individualized counterfactual prediction—without explicitly modeling SCMs or relying on auxiliary information—remain scarce. Although [61, 39] were not originally designed for counterfactuals, we can adapt them by incorporating factual information into intermediate steps and use them as baselines.
Diffusion and Flow Models: Diffusion models [23, 52, 54] and flow-based generative models [10, 58, 36, 4] have demonstrated strong performance across a range of applications. Recently, diffusion and flow-based generative models have been applied to causal inference tasks such as counterfactual image generation [48, 16], causal query answering [9], and inference over causal graphs or structural causal models [8, 29]. To our knowledge, two generative algorithms have been proposed specifically for the potential outcomes (PO) task: the diffusion-based DiffPO [39] and the discrete normalizing flow-based INFs [41]. We conduct extensive comparisons with both methods in this work.
@@ 2.1 Neural ODE and Continuous Normalizing Flow:
$\small\frac{dy(t)}{dt}=v_{\theta}(y(t),t),\quad t\in[0,1],$ (1)
where $v_{\theta}:\mathbb{R}^{d}\times[0,1]\to\mathbb{R}^{d}$ represents the velocity field, which is parameterized by a neural network with parameter $\theta$. The time horizon is rescaled to $[0,1]$ without loss of generality.
@@
$\small\partial_{t}p(y,t)+\nabla\cdot(p(y,t)v_{\theta}(y,t))=0,\quad t\in[0,1],$ (2)
where $\nabla\cdot$ denotes the divergence operator.
When the Neural ODE is well-posed, it defines a continuous and invertible mapping from the initial sample $y_{0}$ to the terminal sample $y_{1}$. The inverse mapping from $y_{1}$ to $y_{0}$ can be computed by integrating the ODE (1) in reverse time.
This framework allows setting $y_{0}\sim p(\cdot,0)$ as the data distribution and $y_{1}\sim p(\cdot,1)$ as the base (noise) distribution, typically chosen as $N(0,I)$. For notational convenience, we define $p_{Y}(\cdot):=p(\cdot,0)$ as the data distribution and $q(\cdot):=p(\cdot,1)$ as the base (noise) distribution. Throughout this paper, we adopt the common choice $q(\cdot)=N(0,I)$.
@@ 2.2 Flow Matching
$\small\mathcal{L}_{\text{FM}}(\theta)=\mathbb{E}_{t\sim\mathcal{U}[0,1],y\sim p(\cdot,t)}\left[\|v_{\theta}(y,t)-u(y,t)\|^{2}\right],$ (3)
@@
$\small\phi(y_{0},y_{1};t)=(1-t)y_{0}+(t+\sigma_{\text{min}}(1-t))y_{1},$ (5)
where $y_{1}\sim N(0,I)$ and $\sigma_{\text{min}}$ is a small positive hyperparameter that ensures $p(\phi,0)\sim N(y_{0},\sigma_{\text{min}}^{2}I)$. When $\sigma_{\text{min}}=0$, this reduces to the strict linear interpolant $\phi=(1-t)y_{0}+ty_{1}$.
Here, $\phi$ serves as a pre-specified reference trajectory that the ODE solution $y(t)$ should follow at time $t$. Under the linear interpolant defined in (5), the corresponding reference velocity field becomes:
$\small\frac{d\phi}{dt}=(1-\sigma_{\text{min}})y_{1}-y_{0}.$ (6)
@@
$\small\mathcal{L}_{\text{CFM}}(\theta)=\ \mathbb{E}_{t\sim\mathcal{U}[0,1],y_{0}\sim p_{Y}(\cdot),y_{1}\sim q(\cdot)}\|v_{\theta}(\phi,t)-\frac{d\phi}{dt}\|^{2}.$ (7)
@@ 3 Problem Settings
Notations and Setups: We consider a setting where each individual is described by covariates $X\in\mathbb{R}^{d_{X}}$, a treatment assignment $A\in\{0,1\}$, and a potential outcome $Y^{(A)}\in\mathbb{R}$ under treatment $A$. The observational dataset is $\mathcal{D}=\{(y^{(a)}_{i},x_{i},a_{i})\}_{i=1}^{n}\sim p_{Y,X,A}(\cdot)$, where $y_{i}^{(a)}$ is the observed potential outcome under treatment $a_{i}$, and $(x_{i},a_{i})\sim p_{X,A}(\cdot)$ are the associated covariates and treatment.
@@
Counterfactual Outcomes: In contrast to potential outcomes, we define the counterfactual outcome as $Y_{\text{cf}}^{(1-a)}:=Y^{(1-a)}\mid X,A=a,Y^{(a)}$, which is the outcome had the individual received the opposite treatment, conditioned on its observed potential (factual) outcome. For convenience, we denote $Y_{\text{cf}}^{(1-a)}$ and $\hat{Y}_{\text{cf}}^{(1-a)}$ as the true and predicted counterfactual outcomes, respectively.
To ensure the identifiability of potential outcomes (in Section 4.1), we adopt the standard assumptions commonly used in the literature [39, 41, 61]:
@@
(1) Consistency: If an individual receives treatment $A=a$, the observed outcome equals the corresponding potential outcome: $Y=Y^{(a)}$. (2) Unconfoundedness: All confounders are observed, i.e., $\{Y^{(0)},Y^{(1)}\}\perp\!\!\!\perp A\mid X$. (3) Overlap: Treatment assignment is probabilistic: $0<P(A=1\mid X=x)<1$, for all $x\in\mathcal{X}$.
In Section 4.2, we will introduce additional assumptions on counterfactual outcomes estimation in PO-Flow and provide the corresponding theoretical analyses for counterfactual recovery.
@@ 4 PO-Flow: A Continuous Normalizing Flow for Potential Outcomes
We introduce PO-Flow, a continuous normalizing flow framework for modeling potential and counterfactual outcomes, with three main goals: 1. predicting individualized potential outcomes, 2. predicting counterfactual outcomes, and 3. learning the densities.
Our PO-Flow framework is built on Continuous Normalizing Flows (CNFs) and is trained using the Flow Matching (FM) objective. Our learning target is the trajectory of the sample $y(t)$ over $t\in[0,1]$. To condition the dynamics on both covariates $x$ and treatment $a$, the Neural ODE is defined as:
$\small\frac{dy(t)}{dt}=v_{\theta}(y(t),t;x,a),\quad t\in[0,1],$ (8)
where $v_{\theta}$ is parametrized by a neural network, with $y(0)$ following the data distribution and $y(1)\sim q(\cdot)$.
@@
$\min_{\theta}\ \mathcal{L}_{\text{CFM}}(\theta)=\ \mathbb{E}_{t\sim\mathcal{U}[0,1],p_{Y,X,A}(\cdot),q(\cdot)}\|v_{\theta}(\phi,t;x,a)-\frac{d\phi}{dt}\|^{2},$ (9)
where $\phi$ and $\frac{d\phi}{dt}$ are defined in (5) and (6), respectively, with $y_{0}=y^{(a)}$ and $y_{1}\sim q(\cdot)$.
Remark: Following prior works [20, 39], one may estimate propensity scores $w(x)=p(A=1\mid X=x)$ with a pre-trained classifier and apply inverse weighting $\frac{1}{w(x)}$ in the loss. We found this provided little performance improvement, but both implementations are included in our code.
Forward Process: We treat the forward process as an encoding operation, denoted by the function $Z:=\Phi_{\theta}(Y^{(A)};\,X,A)$. Given the velocity field $v_{\theta}$, the forward process refers to pushing a factual outcome $y^{(a)}$ (at $t=0$) to a latent embedding $z$ (at $t=1$) via:
$z=\Phi_{\theta}(y^{(a)};\,x,a)=y^{(a)}+\int_{0}^{1}v_{\theta}(y(t),t;x,a)dt,\ \text{with}\ y(0)=y^{(a)}.$ (10)
Through the forward process, $z$ encodes information from the factual outcome $y^{(a)}$ as well as its associated covariates $x$ and treatment $a$.
Reverse Process: We treat the reverse process as a decoding operation. It is initialized differently for potential outcome (Section 4.1) and counterfactual prediction (Section 4.2). In general, at $t=1$, given a latent representation $z$, which can be either sampled from $q(\cdot)=N(0,1)$ or obtained by encoding a factual sample, the reverse process is defined as:
$\hat{y}_{0}=\Phi_{\theta}^{-1}(z;x,a)=z-\int_{0}^{1}v_{\theta}(y(t),t;x,a)dt,\ \text{with}\ y(1)=z.$ (11)
Empirically, we use the Runge-Kutta method (Appendix B.3) for numerical integration.
During training, we construct synthetic paths $\phi$ between $y_{0}$ and $y_{1}\sim q(\cdot)$ using the CFM loss conditioned on $y_{0}$. As shown in Proposition A.1 (Appendix), the gradient of the CFM loss is equivalent to that of the original FM loss. Consequently, the learned velocity field $v_{\theta}$ generates $\hat{y}_{0}$ independent of the specific training samples $y_{0}$, and since $y_{1}\sim q(\cdot)$ is random, $v_{\theta}$ induces a distribution over outcomes conditioned on $(x,a)$.